# Strong time-periodic solutions to a multilayered fluid-structure interaction problem

### v1

@@ 5.1. Estimates of the nonlinear terms
We start by establishing estimates of the maximal regularity space in the lemma below. The assertion of (a) is a consequence of the mixed derivative theorem, see for example [54, Thm. 4.5.10], while the assertion of (b) follows from [1, Thm. III.4.10.2].
[…]
@@
The assertion of (a) is a consequence of Subsection 5.1(b) and $\mathrm{B}_{qp}^{4-\nicefrac{{2}}{{p}}}(\omega)\hookrightarrow\mathrm{C}^{1}(\overline{\omega})$, following from Sobolev embeddings, see for instance [58, Thm. 4.6.1], thanks to the assumptions on $p$ and $q$ that especially imply $\frac{1}{p}+\frac{1}{q}<\frac{3}{2}$, or, equivalently, $14-\frac{2}{p}-\frac{2}{q}>$ 1.
@@
The last preparatory step before tackling the estimates of the nonlinear terms is to show suitable estimates of the transform. These estimates can be derived in a similar way as in [41, Prop. 5.2 and Prop. 6.2] upon employing Subsection 5.1(a) and the relation $a(X)=\frac{\nabla X}{\det(\nabla X)}$.
[…]
@@
    $\|\nabla X\|_{\mathrm{L}^{\infty}(0,T;\mathrm{C}(\overline{\Omega_{\mathrm{f}}}))}+\|\nabla Y\|_{\mathrm{L}^{\infty}(0,T;\mathrm{C}(\overline{\Omega_{\mathrm{f}}}))}+\|a\|_{\mathrm{L}^{\infty}(0,T;\mathrm{C}(\overline{\Omega_{\mathrm{f}}}))}+\|b\|_{\mathrm{L}^{\infty}(0,T;\mathrm{C}(\overline{\Omega_{\mathrm{f}}}))}\leq C,\enspace\text{and}\enspace$
@@
$\left\|b_{\alpha i}\frac{\partial a_{ik}}{\partial x_{j}}\left(X\right)a_{jm}\left(X\right)u_{k}u_{m}\right\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{f}}))}\leq CR^{3},\enspace\text{and}\enspace\$
@@
The following proposition is used to establish the contraction property of the solution map to the linearized problem. It can be proved in the same way as Subsection 5.1. We omit the details here and refer to [41, Prop. 5.3] for a rough outline of the proof.
[…]
@@ 5.2. Proof of existence of the strong time-periodic solution
[…]
@@
$\tilde{u}(0,\cdot)&=\tilde{u}(T,\cdot),&&\enspace\text{in}\enspace\Omega_{\mathrm{f}},\\ \tilde{\eta}_{1}(0,\cdot)=\tilde{\eta}_{1}(T,\cdot),\,\tilde{\eta}_{2}(0,\cdot)&=\tilde{\eta}_{2}(T,\cdot),&&\enspace\text{in}\enspace\omega\\ \tilde{d}_{1}(0,\cdot)=\tilde{d}_{1}(T,\cdot),\,\tilde{d}_{2}(0,\cdot)&=\tilde{d}_{2}(T,\cdot),&&\enspace\text{in}\enspace\Omega_{\mathrm{s}}.\end{aligned}\right.$
@@
$F(u,\pi,\eta_{1},\eta_{2})+f\in\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}),\enspace G(u,\pi,\eta_{1},\eta_{2})+g\in\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\omega))\enspace\text{and}\enspace h\in\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3}).$
@@
On the other hand, for $v_{1}$, $v_{2}\in\overline{\mathbb{B}}_{\mathbb{E}_{1}}(0,R)$, by Subsection 5.1, we find $\|\Phi(v_{1})-\Phi(v_{2})\|_{\mathbb{E}_{1}}\leq CR\cdot\|v_{1}-v_{2}\|_{\mathbb{E}_{1}}$. Thus, choosing $R\coloneqq\min\left\{\frac{1}{2},\frac{1}{4C}\right\}$ and $\varepsilon=\varepsilon(R)\coloneqq\frac{R}{4C}$, we get
$\|\Phi(v)\|_{\mathbb{E}_{1}}\leq\frac{R}{2}\enspace\text{and}\enspace\|\Phi(v_{1})-\Phi(v_{2})\|_{\mathbb{E}_{1}}\leq\frac{1}{2}\cdot\|v_{1}-v_{2}\|_{\mathbb{E}_{1}}.$
In other words, $\Phi$ is a self-map and contraction on $\overline{\mathbb{B}}_{\mathbb{E}_{1}}(0,R)$, and the contraction mapping principle yields the existence of a unique fixed point of $\Phi$. By construction, this shows the existence of a unique solution $v=(u,\pi,\eta_{1},\eta_{2},d_{1},d_{2})\in\overline{\mathbb{B}}_{\mathbb{E}_{1}}(0,R)$ to the transformed interaction problem (3.4). Let us observe that the regularities asserted in Theorem 3.1 are a consequence of the definition of $\mathbb{E}_{1}$.
Finally, in order to obtain the last part of the assertion of the theorem, we invoke the remark after Subsection 5.1 and possibly choose $R>0$ smaller in order to guarantee the validity of (3.2) and then also the diffeomorphism property of $X(t,\cdot)$ for all $t\in[0,T]$. ∎
@@
Acknowledgements This research is supported by the Basque Government through the BERC 2022-2025 program and by the Spanish State Research Agency through BCAM Severo Ochoa excellence accreditation CEX2021-01142-S funded by MICIU/AEI/10.13039/501100011033 and through Grant PID2023-146764NB-I00 funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union. Felix Brandt acknowledges the support by the German National Academy of Sciences Leopoldina through the Leopoldina Fellowship Program with grant number LPDS 2024-07 and the support by DFG project FOR 5528. He also extends his gratitude to the Basque Center for Applied Mathematics for their generous hospitality during his visit, where a portion of this work was completed. Arnab Roy would like to thank the Alexander von Humboldt-Foundation, and the support by Grant RYC2022-036183-I funded by MICIU/AEI/10.13039/501100011033 and by ESF+.

### v2

@@ 6.1. Estimates of the nonlinear terms
We start by establishing estimates of the maximal regularity space in the lemma below. The assertion of (a) is a consequence of the mixed derivative theorem, see for example [61, Thm. 4.5.10], while the assertion of (b) follows from [1, Thm. III.4.10.2].
Lemma 6.1.
@@
The assertion of (a) is a consequence of Subsection 6.1(b) and $\mathrm{B}_{qp}^{4-\nicefrac{{2}}{{p}}}(\omega)\hookrightarrow\mathrm{C}^{1}(\overline{\omega})$, following from Sobolev embeddings, see for instance [65, Thm. 4.6.1], thanks to the assumptions on $p$ and $q$ that especially imply $\frac{1}{p}+\frac{1}{q}<\frac{3}{2}$, or, equivalently, $4-\frac{2}{p}-\frac{2}{q}>1$.
@@
The last preparatory step before tackling the estimates of the nonlinear terms is to show suitable estimates of the transform. These estimates can be derived in a similar way as in [48, Prop. 5.2 and Prop. 6.2] upon employing Subsection 6.1(a) and the relation $a(X)=\frac{\nabla X}{\det(\nabla X)}$.
Lemma 6.4.
@@
    $\|\nabla X\|_{\mathrm{L}^{\infty}(0,T;\mathrm{C}(\overline{\Omega_{\mathrm{f}}}))}+\|\nabla Y\|_{\mathrm{L}^{\infty}(0,T;\mathrm{C}(\overline{\Omega_{\mathrm{f}}}))}+\|a\|_{\mathrm{L}^{\infty}(0,T;\mathrm{C}(\overline{\Omega_{\mathrm{f}}}))}+\|b\|_{\mathrm{L}^{\infty}(0,T;\mathrm{C}(\overline{\Omega_{\mathrm{f}}}))}\leq C,\kern 5.0pt\text{and}\kern 5.0pt$
@@
$\left\|b_{\alpha i}\frac{\partial a_{ik}}{\partial x_{j}}\left(X\right)a_{jm}\left(X\right)u_{k}u_{m}\right\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{f}}))}\leq CR^{3},\kern 5.0pt\text{and}\kern 5.0pt\\$
@@
The following proposition is used to establish the contraction property of the solution map to the linearized problem. It can be proved in the same way as Subsection 6.1. We omit the details here and refer to [48, Prop. 5.3] for a rough outline of the proof.
Proposition $6.$
@@ 6.2. Proof of existence of the strong time-periodic solution
(6.4) $\left\{\begin{aligned} \partial_{t}\tilde{u}-\operatorname{div}\sigma_{\mathrm{f}}(\tilde{u},\tilde{\pi})&=F(u,\pi,\eta_{1},\eta_{2})+f,\,\operatorname{div}\tilde{u}=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\Omega_{\mathrm{f}},\\ \partial_{t}\tilde{\eta}_{1}=\tilde{\eta}_{2},\,\partial_{t}\tilde{\eta}_{2}+P_{\mathrm{m}}\Delta_{s}^{2}\tilde{\eta}_{1}-\Delta_{s}\tilde{\eta}_{2}&=P_{\mathrm{m}}\left(G(u,\pi,\eta_{1},\eta_{2})\right)+P_{\mathrm{m}}g\\ &\qquad-P_{\mathrm{m}}\Bigl(\sigma_{\mathrm{f}}(\tilde{u},\tilde{\pi})\bigr|_{\Gamma_{0}}\mathrm{e}_{3}\cdot\mathrm{e}_{3}\Bigr)\\ &\qquad+P_{\mathrm{m}}\Bigl((\sigma_{\mathrm{s}}(\tilde{d}_{1},\tilde{d}_{2})\mathrm{e}_{3})\bigr|_{\Gamma_{0}}\cdot\mathrm{e}_{3}\Bigr),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\omega,\\ \int_{\omega}\tilde{\eta}_{1}(t,s)\,\mathrm{d}s&=0,&&\kern 5.0pt\text{for}\kern 5.0ptt\in(0,T),\\ \partial_{t}\tilde{d}_{1}=\tilde{d}_{2},\,\partial_{t}\tilde{d}_{2}-\operatorname{div}\sigma_{\mathrm{s}}(\tilde{d}_{1},\tilde{d}_{2})&=h,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\Omega_{\mathrm{s}},\\ \tilde{u}&=P_{\mathrm{m}}(\tilde{\eta}_{2})\mathrm{e}_{3},&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\Gamma_{0},\\ \tilde{d}_{1}&=P_{\mathrm{m}}(\tilde{\eta}_{1})\mathrm{e}_{3},&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\Gamma_{0},\\ \tilde{u}&=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\Gamma_{\mathrm{f}},\\ \tilde{d}_{1}&=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\Gamma_{\mathrm{s}},\\ \tilde{\eta}_{1}=\nabla_{s}\tilde{\eta}_{1}\cdot n_{\omega}&=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\omega,\end{aligned}\right$.
@@
$\left\{\begin{aligned} \tilde{u}(0,\cdot)&=\tilde{u}(T,\cdot),&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{f}},\\ \tilde{\eta}_{1}(0,\cdot)=\tilde{\eta}_{1}(T,\cdot),\,\tilde{\eta}_{2}(0,\cdot)&=\tilde{\eta}_{2}(T,\cdot),&&\kern 5.0pt\text{in}\kern 5.0pt\omega,\\ \tilde{d}_{1}(0,\cdot)=\tilde{d}_{1}(T,\cdot),\,\tilde{d}_{2}(0,\cdot)&=\tilde{d}_{2}(T,\cdot),&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{s}}.\end{aligned}\right$.
@@
$F(u,\pi,\eta_{1},\eta_{2})+f\in\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}),\kern 5.0ptG(u,\pi,\eta_{1},\eta_{2})+g\in\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\omega))\kern 5.0pt\text{and}\kern 5.0pth\in\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3}).$
@@
On the other hand, for $v_{1}$, $v_{2}\in\overline{\mathbb{B}}_{\mathbb{E}_{1}}(0,R)$, by Subsection 6.1, we find
$\|\Phi(v_{1})-\Phi(v_{2})\|_{\mathbb{E}_{1}}\leq CR\cdot\|v_{1}-v_{2}\|_{\mathbb{E}_{1}}.$
Moreover, by the embedding from Subsection 6.1(b) and the Sobolev embedding $\mathrm{B}_{qp}^{4-\nicefrac{{2}}{{p}}}(\omega)\hookrightarrow\mathrm{L}^{\infty}(\omega)$, which holds for all $p$, $q\in(1,\infty)$, see, e.g., [65, Thm. 4.6.1], for $v\in\overline{\mathbb{B}}_{\mathbb{E}_{1}}(0,R)$, there exists a constant $C_{1}>0$ such that
(6.5) $\|\eta(t,\cdot)\|_{\mathrm{L}^{\infty}(\omega)}\leq C_{1}\cdot\|v\|_{\mathbb{E}_{1}}\leq C_{1}R\kern 5.0pt\text{for all}\kern 5.0ptt\in[0,T].$
Thus, choosing $R\coloneqq\min\left\{\frac{1}{2},\frac{1}{4C},\frac{\delta_{0}}{C_{1}}\right\}$, where $\delta_{0}$ is as introduced in (4.3), and $\varepsilon=\varepsilon(R)\coloneqq\frac{R}{4C}$, we get
$\|\Phi(v)\|_{\mathbb{E}_{1}}\leq\frac{R}{2}\kern 5.0pt\text{and}\kern 5.0pt\|\Phi(v_{1})-\Phi(v_{2})\|_{\mathbb{E}_{1}}\leq\frac{1}{2}\cdot\|v_{1}-v_{2}\|_{\mathbb{E}_{1}}.$
In other words, $\Phi$ is a self-map and contraction on $\overline{\mathbb{B}}_{\mathbb{E}_{1}}(0,R)$, and the contraction mapping principle yields the existence of a unique fixed point of $\Phi$. By construction, this shows the existence of a unique solution $v=(u,\pi,\eta_{1},\eta_{2},d_{1},d_{2})\in\overline{\mathbb{B}}_{\mathbb{E}_{1}}(0,R)$ to the transformed interaction problem (4.4). Let us observe that the regularities asserted in Theorem 4.1 are a consequence of the definition of $\mathbb{E}_{1}$. In addition, thanks to $R\leq\frac{\delta_{0}}{C_{1}}$ and (6.5), it follows that the smallness condition (4.2) is satisfied for all $t\in[0,T]$. This yields the diffeomorphism property of $X(t,\cdot)$ for all $t\in[0,T]$, see also Subsection 6.1, and it also implied that no self-intersection of the domain occurs, i.e., the domain does not degenerate. ∎
@@
Acknowledgements. The authors would like to thank the anonymous referees for their careful reading of the manuscript and their valuable comments. This research is supported by the Basque Government through the BERC 2022-2025 program and by the Spanish State Research Agency through BCAM Severo Ochoa excellence accreditation CEX2021-01142-S funded by MICIU/AEI/10.13039/501100011033 and through Grant PID2023-146764NB-I00 funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union. Felix Brandt acknowledges the support by the German National Academy of Sciences Leopoldina through the Leopoldina Fellowship Program with grant number LPDS 2024-07 and the support by DFG project FOR 5528. He also extends his gratitude to the Basque Center for Applied Mathematics for their generous hospitality during his visit, where a portion of this work was completed. Arnab Roy would like to thank the Alexander von Humboldt-Foundation, and the support by Grant RYC2022-036183-I funded by MICIU/AEI/10.13039/501100011033 and by ESF+.